# An additive two-level parallel variant of the DMRG algorithm with coarse-space correction

### v1

@@ 1. Introduction
The basic idea of using a data-sparse tensor format to approximate a high-dimensional optimization problems is certainly not new. A naive implementation of this idea, nevertheless, runs into computational issues since the most basic tensor formats, namely, the canonical polyadic (CP) decomposition and the Tucker decomposition each have their own drawbacks. Indeed, the set of CP-format tensors with rank bounded by $r\in\mathbb{N}$ is not closed, meaning that the best approximation problem in the CP format is ill-posed (see, e.g., [23, Section 3]). While the Tucker format does not suffer from this ill-posedness problem (see, e.g., [23, Section 4] and [22]), the storage cost of an order-$d$ tensor in Tucker format scales exponentially in $d$ so that the curse of dimensionality remains an issue in applications where $d$ is very large. As we explain in more detail in Section 2, the tensor train format offers– in some sense– the best of both worlds by combining the favorable approximability properties enjoyed by the Tucker format with a linear scaling (in $d$) storage cost.
@@
An essential feature of the alternating minimization algorithms listed above is the sequential nature of the lower-order tensor component minimization. Indeed, it is precisely this sequential nature of the local minimization that allows the interpretation of the ALS algorithm as a non-linear Gauss-Seidel relaxation. Thus, in the case of the TT decomposition where an order-$d$ tensor is decomposed in terms of $d$, order-three tensors, we must perform $d$ (or $d-1$ in the case of MALS and two-site DMRG) local minimization and update steps, one-after-the-other in order to complete a global update. Given the difficulty of implementing sequential algorithms of this nature on parallel computing architectures, it is natural to ask if one can develop a non-sequential variant of DMRG in which the local minimization steps are performed independently of each other in parallel, and a global update is obtained by combining all these local updates in a single step. Note that this strategy can be seen as a two-step domain decomposition method applied to the tensor train (or HT) format.
The aim of the present contribution is to develop this line of reasoning by proposing an additive two-level DMRG algorithm that involves independent, local minimization steps, enriched by a second level coarse space correction. Our proposed algorithm, which is directly inspired by additive Schwarz methods from the domain decomposition literature, is particularly amenable to implementation on parallel, distributed architectures since both the local minimization steps and the construction of the coarse space can be performed in parallel. To our knowledge, the present proposal is the first attempt at a non-sequential variant of DMRG with a coarse-space minimization step; the only prior proposal for non-sequential DMRG that we are familiar with is the work of Stoudenmire and White [39], which, nevertheless, does not contain a coarse-space correction.
The remainder of this article is organized as follows. In the forthcoming Section 2.1, we properly introduce the tensor train format and list its basic properties. In Section 2.2, we describe in detail the classical one-site and two-site DMRG algorithms. Subsequently, in Section 3, we state our proposed additive two-level DMRG algorithm with coarse space correction, and in Section 3.1, we describe, in detail, the coarse-step minimization step. Finally, as a first test of the performance of our proposed algorithm, we compute the ground-state (i.e., lowest) energy levels of some strongly-correlated molecular systems from quantum chemistry. These numerical experiments are the subject of Section 4. Further computational aspects of the algorithm including a cost comparison with classical DMRG are given in the supplementary material.
@@ 2.1. Basic Framework
[…]
We now assume that we are given an energy functional $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}$ with the properties that (i) $\mathcal{J}$ is twice continuously differentiable in $U$ and (ii) $\mathcal{J}$ possesses a local minimum in $U$, and we are interested in the solution to the minimization problem
(1) ${E}^{*}:=\underset{\mathbb{U}\in U}{\min}\;\mathcal{J}(\mathbb{U}).$
@@
(2) $\forall\mathbb{U}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}\colon\qquad\mathcal{J}(\mathbb{U})=\frac{\langle\mathbb{U},\mathbb{A}\mathbb{U}\rangle}{\|\mathbb{U}\|^{2}},$
@@
(3) $\forall\mathbb{U}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\colon\qquad\mathcal{J}(\mathbb{U})=\frac{1}{2}\langle\mathbb{U},\mathbb{A}\mathbb{U}\rangle-\langle\mathbb{F},\mathbb{U}\rangle,$
where $\mathbb{A}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ is a symmetric, positive definite operator and $\mathbb{F}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$. In this case, the minimizer of the energy functional (3) corresponds to the solution of the linear operator equation $\mathbb{A}\mathbb{U}=\mathbb{F}$, such equations arising naturally when discretizing high-dimensional, symmetric, elliptic source problems.
@@
Let $\{n_{j}\}_{j=1}^{k}\subset\mathbb{N}$ and let $\mathbb{U}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{k}}$ be a tensor of order $k$. For any $j\in\{1,\ldots,k-1\}$, we write $\mathbb{U}^{<j>}\in\mathbb{R}^{n_{1}\ldots n_{j}\times n_{j+1}\ldots n_{k}}$ to denote the $j^{\rm th}$ canonical unfolding of $\mathbb{U}$. Note that this is a matricization of $\mathbb{U}$ with $\prod_{\ell=1}^{j}n_{j}$ rows and $\prod_{\ell=j+1}^{k}n_{j}$ columns.
@@
Let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$, let $\mathbb{U}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ be a tensor of order $d\in\mathbb{N}$ and for each $j\in\{1,\ldots,d\}$ let $\mathbb{U}^{<j>}\in\mathbb{R}^{n_{1}\ldots n_{j}\times n_{j+1}\ldots n_{d}}$ denote the $j^{\rm th}$ canonical unfolding of $\mathbb{U}$. Then we define the $j^{\rm th}$ separation rank of $\mathbb{U}$, denoted $\text{\rm rank}_{{\rm sep}}(\mathbb{U})\in\mathbb{N}$ as
(4) $\text{\rm rank}_{{\rm sep},j}(\mathbb{U}):=\text{\rm rank}\big{(}\mathbb{U}^{<j>}\big{)}.$
Moreover, we define the separation rank of $\mathbb{U}$, denoted $\text{\rm rank}_{{\rm sep}}(\mathbb{U})\in\mathbb{N}^{d-1}$, as
(5) $\text{\rm rank}_{{\rm sep}}(\mathbb{U}):=\Big{(}\text{\rm rank}_{{\rm sep},1}(\mathbb{U}),\text{\rm rank}_{{\rm sep},2}(\mathbb{U}),\ldots,\text{\rm rank}_{{\rm sep},d-1}(\mathbb{U})\Big{)}.$
@@
Let $\mathbb{r}=(r_{1},\ldots r_{d-1})\subset\mathbb{N}^{d}$ be a $(d-1)$-tuple of natural numbers and let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$. We define the sets $\mathcal{T}_{\leq\mathbb{r}}\subset\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ and $\mathcal{T}_{\mathbb{r}}\subset\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ as
$\mathcal{T}_{\leq\mathbb{r}}:=$$\Big{\{}\mathbb{U}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\colon\qquad\text{\rm rank}_{{\rm sep},j}(\mathbb{U})\leq r_{j}\quad\forall j\in\{1,\ldots,d\}\Big{\}},\quad\text{and}$
$\mathcal{T}_{\mathbb{r}}:=$$\Big{\{}\mathbb{U}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\colon\qquad\text{\rm rank}_{{\rm sep},j}(\mathbb{U})=r_{j}\quad\forall j\in\{1,\ldots,d\}\Big{\}}.$
@@
Returning now to the minimization problem (1), we observe that we can approximate the sought-after minimum $\mathcal{E}^{*}$ by considering, for a judicious choice of ranks $\mathbb{r}\in\mathbb{N}^{d-1}$, the alternative minimization problem
(6) $E^{*}_{\mathbb{r}}:=\underset{\mathbb{U}\in\mathcal{T}_{\mathbb{r}}\cap U}{\min}\;\mathcal{J}(\mathbb{U}).$
Note that elements of $\mathcal{T}_{\mathbb{r}}$ are still order $d$ tensors, and it is, therefore, not clear that the minimization problem (6) is computationally more tractable than the original minimization problem (1). It turns, however, that elements of the manifold $\mathcal{T}_{\mathbb{r}}$ have a data-sparse representation in terms of the so-called tensor train decomposition. To see this, we next introduce the so-called tensor train parameter space.
@@
(7) $\mathcal{U}_{\mathbb{r}}:=\Big{\{}\big{(}\mathbb{X}_{1},\ldots,\mathbb{X}_{d}\big{)}\in\overline{\mathcal{U}}_{\mathbb{r}}\colon\quad\text{\rm rank}(\mathbb{X}_{j}^{<2>})=\text{\rm rank}(\mathbb{X}_{j+1}^{<1>})=r_{j}\quad\forall j=1,\ldots,d-1\Big{\}}.$
Consider Definition 4 of the tensor train parameter space $\overline{\mathcal{U}}_{\mathbb{r}}$ and subset $\mathcal{U}_{\mathbb{r}}$. Three remarks are in order. First, as hinted at in this definition, elements of $\overline{\mathcal{U}}$ will typically be denoted by $\mathbb{X},\mathbb{Y}$, or $\mathbb{Z}$. Second, given $\mathbb{X}=(\mathbb{X}_{1},\mathbb{X}_{2},\ldots\mathbb{X}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$, we notice that each $\mathbb{X}_{j}$ is a third-order tensor if $j\neq 1,d$ and a second order tensor otherwise, and we will frequently use the notation
$\mathbb{X}_{j}(:,x_{j},:)\in\mathbb{R}^{r_{j-1}\times r_{j}},\qquad j\in\{1,\ldots,d\},~{}x_{j}\in\{1,\ldots n_{j}\}$
@@
Let $\mathbb{r}\in\mathbb{N}^{d-1}$, let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$, and let the tensor train parameter space $\overline{\mathcal{U}}_{\mathbb{r}}$ be defined as in Definition 4. We introduce the tensor train contraction mapping $\tau\colon\overline{\mathcal{U}}_{\mathbb{r}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ as the function with the property that for all $\mathbb{X}=(\mathbb{X}_{1},\ldots,\mathbb{X}_{d})$ $\in\overline{\mathcal{U}}_{\mathbb{r}}$ and all $x_{j}\in\{1,\ldots,n_{j}\},~{}j=1,\ldots,d$ it holds that
$\tau(\mathbb{X})(x_{1},\ldots,x_{d})$ $:=\tau(\mathbb{X}_{1},\ldots,\mathbb{X}_{d})(x_{1},\ldots,x_{d})$
$:=\mathbb{X}_{1}(:,x_{1},:)\mathbb{X}_{2}(:,x_{2},:)\ldots\mathbb{X}_{d}(:,x_{d},:)$
$=\sum_{k_{1}=1}^{r_{1}}\sum_{k_{2}=1}^{r_{2}}\ldots\sum_{k_{d-1}=1}^{r_{d-1}}\mathbb{X}_{1}(1,x_{1},k_{1})\mathbb{X}_{2}(k_{1},x_{2},k_{2})\ldots\mathbb{X}_{d}(k_{d-1},x_{d},1).$
@@
Consequently, given any tensor $\tau(\mathbb{X})\in\mathcal{T}_{\leq\mathbb{r}}$, we refer to $\mathbb{X}=(\mathbb{X}_{1},\ldots\mathbb{X}_{d})$ as the tensor train decomposition of $\tau(\mathbb{X})$, and we refer to $\mathbb{X}_{j}$ as the corresponding $j^{\rm th}$ TT core. Additionally, we will use the phrases separation ranks and TT ranks interchangeably.
It has been shown in [42] that the tensor contraction mapping $\tau\colon\mathcal{U}_{\mathbb{r}}\rightarrow\mathcal{T}_{\mathbb{r}}$ is a submersion, i.e., $\tau|_{\mathcal{U}_{\mathbb{r}}}$ is of constant rank given by $\text{dim}\mathcal{T}_{\mathbb{r}}$. This implies, in particular, that $\tau$ is a differentiable mapping that maps open sets to open sets. On the other hand, it is well-known that $\tau$ is not injective. Indeed, given $\mathbb{r}=(r_{1},\ldots,r_{d})\in\mathbb{N}^{d-1}$, let $\mathbb{A}=(\mathbb{A}_{1},\ldots,\mathbb{A}_{d})$ be a tuple of invertible matrices with each $\mathbb{A}_{i}\in\mathbb{R}^{r_{i}\times r_{i}}$, and let $\mathbb{X}\in\overline{\mathcal{U}}_{\mathbb{r}}$. We can then define $\theta_{\mathbb{X}}(\mathbb{A})=(\widehat{\mathbb{X}}_{1},\ldots\widehat{\mathbb{X}}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$ as
$\widehat{\mathbb{X}}_{1}(\mathbb{x}_{j})={\mathbb{X}}_{1}(\mathbb{x}_{j})\mathbb{A}_{1},\quad\widehat{\mathbb{X}}_{2}(\mathbb{x}_{j})=\mathbb{A}_{1}^{-1}{\mathbb{X}}_{1}(\mathbb{x}_{j})\mathbb{A}_{2},\quad\ldots,\quad\widehat{\mathbb{X}}_{d}(\mathbb{x}_{j})=\mathbb{A}_{d-1}^{-1}{\mathbb{X}}_{d}(\mathbb{x}_{j}).$
where $x_{j}\in\{1,\ldots,n_{j}\}$ for $j=1,\ldots,d$. Clearly, we then have $\tau(\mathbb{X})=\tau(\theta_{\mathbb{X}}(\mathbb{A}))$.
@@
Let $\mathbb{r}\in\mathbb{N}^{d-1}$, let $\overline{\mathcal{U}}_{\mathbb{r}}$ be defined according to Definition 4, let the tensor contraction mapping $\tau\colon\overline{\mathcal{U}}_{\mathbb{r}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots\times n_{d}}$ be defined according to Definition 5, and let $\mathbb{X}=(\mathbb{X}_{1},\ldots,\mathbb{X}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$. Then, given $i\in\{1,\ldots,d\}$, we say that $\mathbb{X}$ is an $i$-orthogonal tensor decomposition of $\tau(\mathbb{X})\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots\times n_{d}}$ if for all $j\in\{1,\ldots,i-1\}$ and all $\ell\in\{i+1,\ldots,d\}$ it holds that
$\big{(}\mathbb{X}_{j}^{<2>}\big{)}^{\intercal}\mathbb{X}_{j}^{<2>}=\mathbb{I}\in\mathbb{R}^{r_{j}\times r_{j}}\text{ and }\mathbb{X}_{j}^{<1>}\big{(}\mathbb{X}_{j}^{<1>}\big{)}^{\intercal}=\mathbb{I}\in\mathbb{R}^{r_{j-1}\times r_{j-1}}.$
Additionally, if $\mathbb{X}$ is a 1-orthogonal (respectively, $d$-orthogonal) tensor decomposition of $\tau(X)$, then we say that $\mathbb{X}$ is a right-orthogonal (respectively left-orthogonal) tensor train decomposition of $\tau(X)$.
@@
(9) $\forall\mathbb{X}\in\mathcal{V}_{\mathbb{r}}\colon\qquad\mathcal{j}(\mathbb{X}):=\mathcal{J}\circ\tau(\mathbb{X}).$

### v2

@@ 1. Introduction
The essential idea of using a data-sparse tensor format to approximate high-dimensional optimization problems is certainly not new. A naive implementation of this idea, nevertheless, runs into computational issues since the most basic tensor formats, namely, the canonical polyadic (CP) decomposition and the Tucker decomposition each have their own drawbacks. Indeed, the set of CP-format tensors with rank bounded by $r\in\mathbb{N}$ is not closed, meaning that the best approximation problem in the CP format is ill-posed (see, e.g., [25, Section 3]). While the Tucker format does not suffer from this ill-posedness problem (see, e.g., [25, Section 4] and [24]), the storage cost of an order-$d$ tensor in Tucker format scales exponentially in $d$ so that the curse of dimensionality remains an issue in applications where $d$ is very large. As we explain in more detail in Section 2, the tensor train format offers– in some sense– the best of both worlds by combining the favorable approximability properties enjoyed by the Tucker format with a linear scaling (in $d$) storage cost.
@@
An essential feature of the alternating minimization algorithms listed above is the sequential nature of the lower-order tensor component minimization. Indeed, it is precisely this sequential nature of the local minimization that allows the interpretation of the ALS algorithm as a non-linear Gauss-Seidel relaxation. Thus, in the case of the TT decomposition where an order-$d$ tensor is decomposed in terms of $d$, order-three tensors, we must perform $d$ (or d1d-1 in the case of MALS and two-site DMRG) local minimization and update steps, one-after-the-other in order to complete a global update. Given the difficulty of implementing sequential algorithms of this nature on parallel computing architectures, it is natural to ask if one can develop a non-sequential variant of DMRG in which the local minimization steps are performed independently of each other in parallel, and a global update is obtained by combining all these local updates in a single step. Note that this strategy can be seen as a two-step domain decomposition method applied to the tensor train (or HT) format.
The aim of the present contribution is to develop this line of reasoning by proposing an additive two-level DMRG algorithm that involves independent, local minimization steps, enriched by a second level coarse space correction. Our proposed algorithm, which is directly inspired by additive Schwarz methods from the domain decomposition literature (see Remark 15 below), is well-suited for implementation on distributed architectures since both the local minimization steps and the coarse space construction can be performed in parallel. To our knowledge, the present proposal is the first attempt at a non-sequential variant of DMRG with a coarse-space minimization step; the only prior proposal for non-sequential DMRG that we are familiar with is due to Stoudenmire and White [41], which does not contain coarse-space correction.
The remainder of this article is organized as follows. In the upcoming Sections 2.1 and 2.2, we introduce the tensor train format and the classical one-site and two-site DMRG algorithms respectively. In Section 3, we state our proposed additive two-level DMRG algorithm with coarse space correction, and in Section 3.1, we describe, in detail, the coarse-step minimization step. Finally, as a first test of the performance of our proposed algorithm, we compute the ground-state (lowest) energy levels of some strongly-correlated molecular systems from quantum chemistry. Strongly correlated molecules possess ground state eigenfunctions that are poorly approximated using a single Slater determinants [11] (as in Hartree-Fock or DFT approaches), and are therefore prime candidates for the application of DMRG. These numerical experiments are the subject of Section 4. Further computational aspects of the algorithm including a cost comparison with classical DMRG are given in the supplementary material.
@@ 2.1. Basic Framework
Let $d\in\mathbb{N}$ and $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}^{d}$ be given. We denote by $\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$, the space of order $d$ tensors, and by $U\subset\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ some open set. Following standard convention, elements of the tensor space $\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ will be denoted as $\mathbb{X},\mathbb{Y}$, or $\mathbb{Z}$, etc, and we will assume, throughout this article, that all Euclidean spaces including $\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ are equipped with the Frobenius inner product $\langle\cdot,\cdot\rangle$ and corresponding Frobenius norm $\|\cdot\|$.
We assume that we are given a twice continuously differentiable energy functional $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}$ that possesses a (global) minimum in $U$, and we are interested in the solution to the minimization problem
(1) ${E}^{*}:=\underset{\mathbb{X}\in U}{\min}\;\mathcal{J}(\mathbb{X}).$
@@
(2) $\forall\mathbb{X}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}\colon\qquad\mathcal{J}(\mathbb{X})=\frac{\langle\mathbb{X},\mathbb{A}\mathbb{X}\rangle}{\|\mathbb{X}\|^{2}},$
@@
(3) $\forall\mathbb{X}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\colon\qquad\mathcal{J}(\mathbb{X})=\frac{1}{2}\langle\mathbb{X},\mathbb{A}\mathbb{X}\rangle-\langle\mathbb{F},\mathbb{X}\rangle,$
where $\mathbb{A}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ is a symmetric, positive definite operator and $\mathbb{F}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$. In this case, the minimizer of the energy functional (3) corresponds to the solution of the linear operator equation $\mathbb{A}\mathbb{X}=\mathbb{F}$, such equations arising naturally when discretizing high-dimensional, symmetric, elliptic source problems.
@@
Let $\{n_{j}\}_{j=1}^{k}\subset\mathbb{N}$ and let $\mathbb{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{k}}$ be a tensor of order $k$. For any $j\in\{1,\ldots,k-1\}$, we write $\mathbb{X}^{<j>}\in\mathbb{R}^{n_{1}\ldots n_{j}\times n_{j+1}\ldots n_{k}}$ to denote the $j^{\rm th}$ canonical unfolding of $\mathbb{U}$. Note that this is a matricization of $\mathbb{X}$ with $\prod_{\ell=1}^{j}n_{j}$ rows and $\prod_{\ell=j+1}^{k}n_{j}$ columns. By convention, $\mathbb{X}^{<d>}\in\mathbb{R}^{n_{1}\ldots n_{d}\times 1}$ denotes the vectorization of $\mathbb{U}$.
@@
Let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$, let $\mathbb{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ be a tensor of order $d\in\mathbb{N}$ and for each $j\in\{1,\ldots,d\}$ let $\mathbb{X}^{<j>}\in\mathbb{R}^{n_{1}\ldots n_{j}\times n_{j+1}\ldots n_{d}}$ denote the $j^{\rm th}$ canonical unfolding of $\mathbb{X}$. Then we define the $j^{\rm th}$ separation rank of $\mathbb{X}$, denoted $\text{\rm rank}_{{\rm sep}}(\mathbb{X})\in\mathbb{N}$ as
(4) $\text{\rm rank}_{{\rm sep},j}(\mathbb{X}):=\text{\rm rank}\big(\mathbb{X}^{<j>}\big).$
Moreover, we define the separation rank of $\mathbb{X}$, denoted $\text{\rm rank}_{{\rm sep}}(\mathbb{X})\in\mathbb{N}^{d-1}$, as
(5) $\text{\rm rank}_{{\rm sep}}(\mathbb{X}):=\Big(\text{\rm rank}_{{\rm sep},1}(\mathbb{X}),\text{\rm rank}_{{\rm sep},2}(\mathbb{X}),\ldots,\text{\rm rank}_{{\rm sep},d-1}(\mathbb{X})\Big).$
@@
Let $\mathbb{r}=(r_{1},\ldots r_{d-1})\subset\mathbb{N}^{d}$ be a (d1)(d-1)-tuple of natural numbers and let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$. We define the sets $\mathcal{T}_{\leq\mathbb{r}}\subset\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ and $\mathcal{T}_{\mathbb{r}}\subset\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ as
$\mathcal{T}_{\leq\mathbb{r}}:=$ $\Big\{\mathbb{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\colon\qquad\text{\rm rank}_{{\rm sep},j}(\mathbb{X})\leq r_{j}\quad\forall j\in\{1,\ldots,d\}\Big\},\quad\text{and}$
$\mathcal{T}_{\mathbb{r}}:=$ $\Big\{\mathbb{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\colon\qquad\text{\rm rank}_{{\rm sep},j}(\mathbb{X})=r_{j}\quad\forall j\in\{1,\ldots,d\}\Big\}.$
@@
Returning now to the minimization problem (1), we observe that we can approximate the sought minimum $\mathcal{E}^{*}$ by considering, for a judicious choice of ranks $\mathbb{r}\in\mathbb{N}^{d-1}$, the alternative minimization problem
(6) $E^{*}_{\mathbb{r}}:=\underset{\mathbb{X}\in\mathcal{T}_{\mathbb{r}}\cap U}{\min}\;\mathcal{J}(\mathbb{X}).$
Note that elements of $\mathcal{T}_{\mathbb{r}}$ are still order $d$ tensors, and it is, therefore, not clear that the minimization problem (6) is computationally more tractable than the original minimization problem (1). It turns out, however, that elements of the manifold $\mathcal{T}_{\mathbb{r}}$ have a data-sparse representation in terms of the so-called tensor train decomposition. To see this, we next introduce the so-called tensor train parameter space.
@@
(7) $\mathcal{U}_{\mathbb{r}}:=\Big\{\big(\mathbb{U}_{1},\ldots,\mathbb{U}_{d}\big)\in\overline{\mathcal{U}}_{\mathbb{r}}\colon\quad\text{\rm rank}(\mathbb{U}_{j}^{<2>})=\text{\rm rank}(\mathbb{U}_{j+1}^{<1>})=r_{j}\quad\forall j=1,\ldots,d-1\Big\}.$
Consider Definition 4 of the tensor train parameter space $\overline{\mathcal{U}}_{\mathbb{r}}$ and subset $\mathcal{U}_{\mathbb{r}}$. Three remarks are in order. First, as hinted at in this definition, elements of $\overline{\mathcal{U}}$ will typically be denoted by $\mathbb{U},\mathbb{V}$, or $\mathbb{W}$. Second, given $\mathbb{U}=(\mathbb{U}_{1},\mathbb{U}_{2},\ldots\mathbb{U}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$, we notice that each $\mathbb{U}_{j}$ is a third-order tensor if $j\neq 1,d$ and a second order tensor otherwise, and we will frequently use the notation
$\mathbb{U}_{j}(:,x_{j},:)\in\mathbb{R}^{r_{j-1}\times r_{j}},\qquad j\in\{1,\ldots,d\},~x_{j}\in\{1,\ldots n_{j}\}$
@@
Let $\mathbb{r}\in\mathbb{N}^{d-1}$, let $\{n_{j}\}_{j=1}^{d}\subset\mathbb{N}$, and let the tensor train parameter space $\overline{\mathcal{U}}_{\mathbb{r}}$ be defined as in Definition 4. We introduce the tensor train contraction mapping $\tau\colon\overline{\mathcal{U}}_{\mathbb{r}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ as the function with the property that for all $\mathbb{U}=(\mathbb{U}_{1},\ldots,\mathbb{U}_{d})$ $\in\overline{\mathcal{U}}_{\mathbb{r}}$ and all $x_{j}\in\{1,\ldots,n_{j}\},~j=1,\ldots,d$ it holds that
$\tau(\mathbb{U})(x_{1},\ldots,x_{d})$ $:=\tau(\mathbb{U}_{1},\ldots,\mathbb{U}_{d})(x_{1},\ldots,x_{d})$
$:=\mathbb{U}_{1}(:,x_{1},:)\mathbb{X}_{2}(:,x_{2},:)\ldots\mathbb{U}_{d}(:,x_{d},:)$
$=\sum_{k_{1}=1}^{r_{1}}\sum_{k_{2}=1}^{r_{2}}\ldots\sum_{k_{d-1}=1}^{r_{d-1}}\mathbb{U}_{1}(1,x_{1},k_{1})\mathbb{U}_{2}(k_{1},x_{2},k_{2})\ldots\mathbb{U}_{d}(k_{d-1},x_{d},1).$
@@
Consequently, given any tensor $\tau(\mathbb{U})\in\mathcal{T}_{\leq\mathbb{r}}$, we refer to $\mathbb{U}=(\mathbb{U}_{1},\ldots\mathbb{U}_{d})$ as a tensor train decomposition of $\tau(\mathbb{U})$, and we refer to $\mathbb{U}_{j}$ as the corresponding $j^{\rm th}$ TT core. Additionally, we will use the phrases separation ranks and TT ranks interchangeably.
It has been shown in [44] that the tensor contraction mapping $\tau\colon\mathcal{U}_{\mathbb{r}}\rightarrow\mathcal{T}_{\mathbb{r}}$ is a submersion, i.e., $\tau|_{\mathcal{U}_{\mathbb{r}}}$ is of constant rank given by $\text{dim}\mathcal{T}_{\mathbb{r}}$. This implies, in particular, that $\tau$ is a differentiable mapping that maps open sets to open sets. On the other hand, it is well-known that $\tau$ is not injective. Indeed, given $\mathbb{r}=(r_{1},\ldots,r_{d})\in\mathbb{N}^{d-1}$, let $\mathbb{A}=(\mathbb{A}_{1},\ldots,\mathbb{A}_{d})$ be a tuple of invertible matrices with each $\mathbb{A}_{i}\in\mathbb{R}^{r_{i}\times r_{i}}$, and let $\mathbb{U}\in\overline{\mathcal{U}}_{\mathbb{r}}$. We can then define $\theta_{\mathbb{U}}(\mathbb{A})=(\widehat{\mathbb{U}}_{1},\ldots\widehat{\mathbb{U}}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$ as
$\widehat{\mathbb{U}}_{1}({x}_{1})={\mathbb{U}}_{1}({x}_{1})\mathbb{A}_{1},\quad\widehat{\mathbb{U}}_{2}({x}_{2})=\mathbb{A}_{1}^{-1}{\mathbb{U}}_{2}({x}_{2})\mathbb{A}_{2},\quad\ldots,\quad\widehat{\mathbb{U}}_{d}({x}_{d})=\mathbb{A}_{d-1}^{-1}{\mathbb{X}}_{d}({x}_{d}).$
where each $x_{j}\in\{1,\ldots,n_{j}\}$ for $j=1,\ldots,d$. Clearly, we then have $\tau(\mathbb{U})=\tau(\theta_{\mathbb{U}}(\mathbb{A}))$.
@@
Let $\mathbb{r}\in\mathbb{N}^{d-1}$, let $\overline{\mathcal{U}}_{\mathbb{r}}$ be defined according to Definition 4, let the tensor contraction mapping $\tau\colon\overline{\mathcal{U}}_{\mathbb{r}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots\times n_{d}}$ be defined according to Definition 5, and let $\mathbb{U}=(\mathbb{U}_{1},\ldots,\mathbb{U}_{d})\in\overline{\mathcal{U}}_{\mathbb{r}}$. Then, given $i\in\{1,\ldots,d\}$, we say that $\mathbb{U}$ is an $i$-orthogonal tensor decomposition of $\tau(\mathbb{U})\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots\times n_{d}}$ if and only if for all $j\in\{1,\ldots,i-1\}$ and all $\ell\in\{i+1,\ldots,d\}$ it holds that
$\big(\mathbb{U}_{j}^{<2>}\big)^{\intercal}\mathbb{U}_{j}^{<2>}=\mathbb{I}\in\mathbb{R}^{r_{j}\times r_{j}}\text{ and }\mathbb{U}_{\ell}^{<1>}\big(\mathbb{U}_{\ell}^{<1>}\big)^{\intercal}=\mathbb{I}\in\mathbb{R}^{r_{\ell-1}\times r_{\ell-1}}.$
Additionally, if $\mathbb{U}$ is a 1-orthogonal (respectively, $d$-orthogonal) tensor decomposition of $\tau(\mathbb{U})$, then we say that $\mathbb{U}$ is a right-orthogonal (respectively left-orthogonal) tensor train decomposition of $\tau(\mathbb{U})$.
@@
(9) $\forall\mathbb{U}\in\mathcal{V}_{\mathbb{r}}\colon\qquad\mathcal{j}(\mathbb{U}):=\mathcal{J}\circ\tau(\mathbb{U}).$